Dep.= % WRK Indep.= % MGT
SUMMARY OUTPUT
Regression Statistics
Multiple R
R Square
Adjusted R Square
Standard Error
Observations
ANOVA
Significance
df SS MS F F
Regression 102.1488 148.9539
Residual
Total 12.0000
Standard
Coefficients Error t Stat P-value Lower 95% Upper 95%
Intercept
% MGT 0.4543
SE CI CI PI PI
Predicted Predicted Lower Upper Lower Upper
x0 Value Value 95% 95% 95% 95%
67.0000 67.8474 65.8779 69.8169
72.0000 70.1189 68.2003 72.0375
76.0000 71.9361 69.7884 74.0838
Dep.= % MGT Indep.= % WRK
SUMMARY OUTPUT
Regression Statistics
Multiple R
R Square
Adjusted R Square
Standard Error
Observations
ANOVA
Significance
df SS MS F F
Regression 460.8873 148.9539
Residual
Total 12.0000
Standard
Coefficients Error t Stat P-value Lower 95% Upper 95%
Intercept
% WRK 2.0498
SE CI CI PI PI
Predicted Predicted Lower Upper Lower Upper
x0 Value Value 95% 95% 95% 95%
67.0000 65.6122 61.4287 69.7957
72.0000 75.8612 71.7859 79.9365
76.0000 84.0603 79.4982 88.6224

Is there a relationship between the percent of management jobs held by men in a company (% MGT) and the percent of men working in a company (% WRK)? Based on this paragraph of text, use the correct excel output above to answer the following question.

The Least Squares Method determines the values of b0 and b1 that minimize this sum. What is the resulting value?

a.

0.6858

b.

3.0942

c.

None of the answers is correct

d.

7.5435

e.

34.0358
Answer #1

The Least Squares Method determines the values of b0 and b1 that minimize this sum which is sum of square of error(residuals)

from above: SS(error) =df(residual)*MS(regression/F) =11*102.1488/148.9539 =7.5435

df SS MS F
regression 1 102.1488 102.1488 148.9539
residual 11 7.543521 0.685775
total 12

option D is correct : 7.5435